OpenStudy (anonymous):

Solve the following equation in the real number system. Please show all of your work. x^3+11/2x^2+5x-4=0

OpenStudy (anonymous):

x^3+(11/2)x^2+5x-4=0 (x+2)(x+4)(x-1/2)=0 x=1/2, x=-2,x=-4 ans.....

OpenStudy (anonymous):

Thanx!

OpenStudy (anonymous):

your wc

OpenStudy (anonymous):

here you can start by dividing the whole prob with trial and errors like x=-2,-4, etc until you get them all...good luck diah...